40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 16 17:26:23 2010, Jenne, Update, PEM, Guralp Breakout Box pulled 
    Reply  Tue Apr 20 00:32:30 2010, Jenne, Update, PEM, Guralp Breakout Box put back Gur2X_before_20Apr2010.jpgGur2X_after_replace_100k_resistors_20Apr2010.jpg
Message ID: 2811     Entry time: Tue Apr 20 00:32:30 2010     In reply to: 2802
Author: Jenne 
Type: Update 
Category: PEM 
Subject: Guralp Breakout Box put back 

Quote:

I pulled the Guralp breakout box from the rack, and it's sitting on the EE bench here.  The game plan is to check out the Gur2X channel.  

Rana and Steve have been investigating, and found that the X channel has been funky (which has been known for ~a month or two) when the seismometer has been plugged in, and also when the seismometers have been unplugged, but the box is left on.  The funkyness goes away when the box is turned off.  Since it's not there when the box is off, it seems that it's not a problem with the cable from the box to the ADC, or in the ADC channel.  Since it is there when the box is on, but the seismometer is unplugged, it's clear that it's probably in the box itself.  

Preliminarily, I've connected a set of BNC clipdoodles to the input testpoints, and another set to the output.  They're both connected to a 'scope (which is on it's battery so it's not connected to any Ground), and when I tap on the circuit board the input trace is totally unchanged, but the output trace goes kind of crazy, and gets more fuzzy, and picks up a DC offset.  Koji is concerned that some of the big capacitors may have an iffy connection to the board. 

Investigations will continue Monday morning. 

 The Guralp Box appears to be back in working order.  It's reinstalled with the 2 seismometers plugged in.

In order:

* Koji suggested retouching the through-board solder joints on the broken channel (EW2 = Gur2X) with a bit of solder to ensure the connections were good.  Check.

* "C7", one of the giant 1uF capacitors on each channel is totally bypassed, and since that was one of the original suspects, Rana removed the (possibly) offending capacitor from EW2.

* Rana and I isolated the craziness to the final differential output stage.  We tried each of the testpoints after the individual gain / filter stages, and found that the signals were all fine, until after the output stage.

* I started to remove the resistors in the output stage (with the plan to go through the resistors, capacitors, and even the amplifier chip if neccessary), and noticed that 2 of the 1k resistors came off too easily, as if they were just barely connected in the first place.  After replacing only the 4 1k resistors, the craziness seemed to be gone.  I poked and gently bent the board, but the output wouldn't go crazy.  I declared victory.

* I reinstalled the box in its normal spot, and put Gur2 (which had been out by the bench for use as a test signal) back next to the other seismometers.  We are in nominal condition, and should be able to do a huddle test this week.

 

I looked at the time traces of all the seismometer channels, and they all looked good.  I'll put a spectra in in the morning....I'm too impatient to wait around for the low frequency FFTs.

 

Attached are the before and after pictures of the output stage of EW2 / Gur2X.  The "before" is the one with the OUT+ and OUT- words upsidedown.  The "after" picture has them right side up.

Attachment 1: Gur2X_before_20Apr2010.jpg  68 kB  Uploaded Tue Apr 20 01:33:30 2010  | Hide | Hide all
Gur2X_before_20Apr2010.jpg
Attachment 2: Gur2X_after_replace_100k_resistors_20Apr2010.jpg  84 kB  Uploaded Tue Apr 20 01:33:44 2010  | Hide | Hide all
Gur2X_after_replace_100k_resistors_20Apr2010.jpg
ELOG V3.1.3-